Calculate the mass of magnesium that would be deposited by 8 faradays of electricity
(Mg - 24)
1
Expert's answer
2020-03-11T09:06:49-0400

m = (QM)/(zF) = (8·96485·24)/(2·96485) = 96 g

Answer: 96 g.